# Homotopy Methods for Convex Optimization

### v1

@@ 3.1.3 A Regular Homotopy for Real Zero Polynomials
Let $p_{0},p_{1}$ be smooth real zero polynomials with bounded rigidly convex sets $\mathcal{R}(p_{0})$ and $\mathcal{R}(p_{1})$. Moreover, let $\varepsilon\colon[0,1]\to[0,\infty)$ be a smooth function with $\varepsilon(0)=\varepsilon(1)=0$ and $\varepsilon(t)>0$ for $t\in(0,1)$. We define the homotopies
@@
    $p_{t}(\mathbf{x})\coloneqq\mathcal{F}_{\varepsilon(t)}[\hat{p}_{t}](\mathbf{x}).$
Proposition 14.
@@
    If $p_{0}$ and $p_{1}$ are smooth, then $p_{t}$ is smooth.
@@
(ii) If $p_{0}$ and $p_{1}$ are smooth, we have that $s\mapsto\hat{p}_{t}(s\mathbf{x})$ has zeros of multiplicity at most $2,$ for every $t\in(0,1)$ and $0\neq\mathbf{x}\in\mathbb{R}^{n}$. Hence, applying the smoothing operator $\mathcal{F}_{\varepsilon(t)}$ gives rise to a smooth polynomial $p_{t}$ for every $t\in[0,1]$. ∎
Note that $\hat{p}_{t}$ is in general not smooth, even if $p_{0}$ and $p_{1}$ are smooth (see for example Fig. 2 for $\varepsilon=0$ and $t=0.6$)
@@
If $p_{0},p_{1}$ are smooth and $\mathcal{R}(p_{0}),\mathcal{R}(p_{1})$ are bounded, then the homotopy $p_{t}$ from Definition 13 is regular, and therefore the induced differential equation (6) has a unique solution.
@@
Regularity follows from smoothness of the $p_{t}$. Indeed, if a point leaves the innermost ring of zeros of $p_{t}$ over time, this would imply a multiple zero of some $p_{t}$ on some line. ∎
Remark 16.
A similar statement to Proposition 14 can also be shown when dropping the smoothness-assumption for $\mathcal{R}(p_{1})$. In this situation $s\mapsto\hat{p}_{t}(s\mathbf{a})$ can have zeros up to multiplicity $d\coloneqq\deg(p_{1})$. Applying $\mathcal{F}_{\varepsilon(t)}$ $d$ times to $\hat{p},$ instead of only once, gives rise to a homotopy that satisfies the assumptions.
Figure 3: Optimizing a linear function on a random spectrahedron.
@@ 4.1 Hyperbolic Programming
For this purpose, let us recall the notion of hyperbolic polynomials. Let $h\colon\mathbb{R}^{n+1}\to\mathbb{R}$ be a homogeneous polynomial. We call $h$ hyperbolic in the direction $\mathbf{e}\in\mathbb{R}^{n+1}$ if the univariate polynomial $t\mapsto h(\mathbf{y}-t\mathbf{e})$ has only real roots, for every $\mathbf{y}\in\mathbb{R}^{n+1}$. Hyperbolic polynomials give rise to a closed convex cone
@@
where $\mathbf{y}=(-\sum_{i=1}^{n}x_{i},x_{1},\ldots,x_{n})$. Moreover, the rigidly convex set $\mathcal{R}(p_{k})$ is bounded for $k\geqslant 2,$ see Appendix C. Note that $\mathcal{R}(p_{k})$ corresponds (up to an isometry) to the set
@@
Regarding (ii), we use the best known spectrahedral representation of $\mathcal{C}_{\mathbf{e}}(p_{k})$ introduced in [7]. According to the construction, the matrix size of this represenation corresponds to the number of strings of length at most $k-1$ with distinct characters from an alphabet of size $n+1$. In other words, the matrix size $s(k,n)$ is given by
@@
$n=3$
$n=4$
$n=5$
$n=6$
$n=7$
$n=8$
$k=n-1$
$5$
$26$
$157$
$\mathbf{1100}$
$\mathbf{8801}$
$\mathbf{79210}$
$k=\lfloor\frac{n+1}{2}\rfloor$
$5$
$6$
$37$
$50$
$\mathbf{401}$
$\mathbf{5861}$
@@
4.2 Optimizing over the $k$-Ellipsoid
@@
Note that $p$ is not differentiable at the focal points $\mathbf{u}_{i}$. Thus even if $\mathbf{u}_{i}\notin\partial\mathcal{E}_{n}(\mathbf{u}_{1},\ldots,\mathbf{u}_{k},r)$, there are times $t$ where the functions $p_{t}(\mathbf{x})$, constructed from the homotopy in Section 3.2, might not be differentiable and strictly concave. However, this happens for at most $k$ many different times. Therefore, the numerical integrator evaluates $p_{t}(\mathbf{x})$ on these problematic points with probability 00. At all other points and times, the function is concave in our above sense, so the method will work with probability $1$. The homotopy method for a $3$-ellipse is shown in Fig. 6.
@@
Note that (16) is a linear matrix inequality whose matrix size is exponential in the number of focal points $k$. There is also no significantly more efficient way of defining the $k$-ellipse by a linear matrix inequality, since the degree of the defining polynomial also increases exponentially with $k$. More precisely, this degree is $2^{k}$ if $k$ is odd and $2^{k}-{k\choose k/2}$ if $k$ is even [25]. Fig. 6 shows the zero set of the real zero polynomial $\det(\mathcal{M}(\mathbf{x}))$ for a $3$-ellipse.
[…]
@@
$k=2$
$k=3$
$k=4$
$k=5$
$k=10$
$k=100$
$n=2$ Homotopy
$0.02(1)$
$0.02(1)$
$0.06(4)$
$0.04(1)$
$0.09(4)$
$1.5(4)$
SDP
$0.03(1)$
$0.07(1)$
$3.3(4)$
$\approx 1500$*
$2^{\text{nd}}$ order cone
$0.02(1)$
$0.02(1)$
$0.03(1)$
$0.03(2)$
$0.04(2)$
$0.15(3)$
$n=4$ Homotopy
$0.05(2)$
$0.05(1)$
$0.07(3)$
$0.10(2)$
$0.21(5)$
$4(1)$
SDP
$0.06(1)$
$23(2)$
$2^{\text{nd}}$ order cone
$0.02(1)$
$0.03(1)$
$0.03(1)$
$0.03(1)$
$0.05(2)$
$0.22(3)$
@@ 4.3 Optimizing over a $p$-Norm Ball
For the SDP representation, note that $B_{8,r}$ is not a spectrahedron (see [23, Example 2.29] for a similar argument for $B_{4,r}$). However, it can be represented as the projection of a spectrahedron. This leads to the following SDP which is equivalent to (18) for $p=8$.
@@ 4.4 Geometric Programming with a Single Constraint
[…]
@@ 5 Conclusion and Open Questions
Moreover, we leave an exact runtime analysis of this method for further work. This would include for example a comparision of the complexity of our method with interior-point methods applied to SDPs. Finally, since there are multiple canonical choices of homotopies for certain optimization classes (see Appendix A and Section 3.1 in the case of SDPs), this suggests to study whether there exist alternative choices of homotopies that lead to computationally more efficient results.
@@ Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [doi:10.55776/P33122]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. AK further acknowledges funding of the Austrian Academy of Sciences (ÖAW) through the DOC scholarship 26547. Moreover, AK wants to thank Hamza Fawzi, and Oisín Faust for fruitful discussions.
Appendix A A Homotopy within Spectrahedra not Satisfying the Assumptions
In the following, we present a canonical homotopy within the class of spectrahedra, which does not satisfy the conditions in Theorem 6. We nevertheless include this homotopy in this paper, since it might be the most intuitive choice for a homotopy to solve semidefinite programs, and it is interesting why this does not work out in general.
Let $A=(A_{1},\ldots,A_{n})$ be a tuple of symmetric $s\times s$ matrices. We define the spectrahedron generated by $A$ as
$\mathcal{S}(A)\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid A(\mathbf{x})=I_{s}+x_{1}A_{1}+\cdots+x_{n}A_{n}\succcurlyeq 0\}$
Note that all those spectrahedra $\mathcal{S}(A)$ contain $\mathbf{0}$ in their interior.
For $p_{A}(\mathbf{x})\coloneqq\det(A(\mathbf{x}))$ we have
$\mathcal{S}(A)=\{\mathbf{x}\in\mathbb{R}^{n}\mid s\mapsto p_{A}(s\mathbf{x})\textrm{ has no roots in }[0,1)\},$
and we have that $\mathbf{x}\in\partial\mathcal{S}(A)$ implies that $p_{A}(\mathbf{x})=0$.
Now let $\mathcal{S}(A),\mathcal{S}(B)$ be two spectrahedra defined by $n$-tuples of $s\times s$ matrices $A$ and $B$. We define the homotopy of matrix equations
$M(\mathbf{x},t)\coloneqq(1-t)A(\mathbf{x})+tB(\mathbf{x}).$
Setting $p_{t}(\mathbf{x})\coloneqq\det(M(\mathbf{x},t))$, we obtain a family of spectrahedra
$\mathcal{S}_{t}\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}\mid s\mapsto p_{t}(s\mathbf{x})\textrm{ has no roots in }[0,1)\}=\mathcal{S}(M(\cdot,t)).$
However, this family of sets does not satisfy the boundedness condition necessary for Theorem 6, even if both $\mathcal{S}(A)$ and $\mathcal{S}(B)$ are bounded, as the following example shows.
Example 19.
Let
$A(\mathbf{x})=\begin{pmatrix}1-x_{1}&x_{2}\\ x_{2}&1+x_{1}\end{pmatrix}\quad\text{ and }B(\mathbf{x})=\begin{pmatrix}1-x_{2}&x_{1}\\ x_{1}&1+x_{2}\end{pmatrix}$
We have that both $p_{A}(\mathbf{x})=p_{B}(\mathbf{x})=1-x_{1}^{2}-x_{2}^{2}$, and thus
$\mathcal{S}(A)=\mathcal{S}(B)=\{\mathbf{x}\in\mathbb{R}^{2}\mid 1-x_{1}^{2}-x_{2}^{2}\geqslant 0\}$
are both the unit disk. Note that we have
$M\left(\mathbf{x},\frac{1}{2}\right)=\begin{pmatrix}1-\frac{x_{1}+x_{2}}{2}&\frac{x_{1}+x_{2}}{2}\\ \frac{x_{1}+x_{2}}{2}&1+\frac{x_{1}+x_{2}}{2}\end{pmatrix}$
and hence
$\mathcal{S}_{\frac{1}{2}}=\big{\{}\mathbf{x}\in\mathbb{R}^{2}\mid|x_{1}+x_{2}|\leqslant\sqrt{2}\big{\}},$
which is an unbounded diagonal strip. We refer to Fig. 9 for a visualization of this homotopy.
Figure 9: A visualization of the spectrahedra arising from the homotopy of matrix equations in Example 19. Since the resulting spectrahedra are not bounded, this is not a valid homotopy for our method.
Appendix B Non-Linear Objective Function and a Homotopy of Objective Functions
In the following, we generalize the homotopy method in two directions:
  • $\triangleright$
    We allow for concave objective functions instead of only linear ones.
  • $\triangleright$
    We allow for a homotopy of objective functions.
More precisely, we set up a differential equation whose solution $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n}$ is the optimal value of
$\max_{\mathbf{x}\in C_{t}}\ f_{t}(\mathbf{x})$ (21)
for every $t\in[0,1],$ if the initial value $\mathbf{a}_{0}$ is the optimal value of the optimization problem for $t=0$.
We again assume that $C_{t}$ is a compact convex set for every $t\in[0,1]$, and we assume to have a regular homotopy $p$ of smooth descriptions, just as before. Moreover, we assume that $(t,\mathbf{x})\mapsto f_{t}(\mathbf{x})$ is three times continuously differentiable, that $f_{t}$ is concave, and $\nabla f_{t}\neq 0$ on $C_{t}$, i.e. all stationary points are outside of $C_{t}$. This guarantees that the optimal solution $\mathbf{a}_{t}$ satisfies $\mathbf{a}_{t}\in\partial C_{t}$.
Using Lagrange multipliers, one analogously obtains that $\mathbf{a}_{t}$ is a stationary point of the constraint problem at time $t$ if and only if
  1. ($\text{C}_{1}$)
    $\nabla f_{t}(\mathbf{a}_{t})=\lambda_{t}\nabla p_{t}(\mathbf{a}_{t})$
  2. ($\text{C}_{2}$)
    $p_{t}(\mathbf{a}_{t})=0$
where $\lambda_{t}\in\mathbb{R}$ is the Lagrange multiplier for the constraint $p_{t}$. If $f_{t}$ has its global maximum outside of $C_{t}$, we have that $\lambda_{t}<0$ (see the proof of Lemma 20).
We translate these two conditions again into an initial value problem, whose solution is a path $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n};t\mapsto\mathbf{a}_{t}$ where $\mathbf{a}_{t}$ is a stationary point of the constrained optimization problem at time $t$.
Differentiating (5) and condition ($\text{C}_{2}$) using Lemma 4, we obtain again the initial value problem for stationary points $\mathbf{a}\colon[0,1]\to\mathbb{R}^{n}$:
where $K$ and $m$ are defined in Eq. 7. Again, the parameter $k\in\{1,\ldots,n\}$ in (7) is chosen such that $\partial_{k}f(\mathbf{a}_{t})\neq 0$ for the optimal path $\mathbf{a}_{t}$. [^7] Depending on the objective function, such a choice of $k$ might not be obvious, and one might change the choice during the time evolution.
In summary, for every solution $\mathbf{a}\colon t\mapsto\mathbf{a}_{t}$, we have that $\mathbf{a}_{t}$ is a stationary point of the constraint optimization problem of $p_{t}$ and $f_{t}$.
We now prove that there is a unique solution for the optimization problem, if $p_{t}$ and $f_{t}$ are (strictly) concave functions. For this purpose, we show that $K(\mathbf{a}_{t},t)$ is invertible. The uniqueness then follows as in Theorem 6.
Lemma 20.
Let $\mathbf{a}_{t}$ be a stationary point of the constrained problem at time $t$. Assume the following:
  1. (i)
    $p_{t}$ and $f_{t}$ are concave on $C_{t}$ where one of them is strict. [^8] The statement can also be relaxed to strict quasi-concave functions instead of strict concave functions.
  2. (ii)
    $\nabla f_{t}\neq 0$ on $C_{t}$
Then $K(\mathbf{a}_{t},t)$ is invertible.
Proof.
We start with an observation necessary for the later argument. Since $\nabla f_{t}\neq 0$ on $C_{t}$, we have that $\mathbf{a}_{t}\in\partial C_{t}$ is an optimal point of the optimization problem on $C_{t}$. Hence, by concavity of $f_{t}$, we have
$\nabla f_{t}(\mathbf{a}_{t})^{t}(\mathbf{x}-\mathbf{a}_{t})\leqslant 0$
We now show that for every $\mathbf{v}\in\mathbb{R}^{n}$, we have that
Assume that $\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0$. This implies $\mathbf{v}^{t}\nabla f_{t}(\mathbf{a}_{t})=0$. Moreover, we have (since $\lambda_{t}<0$) and by (strict) concavity that $\mathbf{v}^{t}A(\mathbf{a}_{t},t)\mathbf{v}<0$, where
For every $i\in\{1,\ldots,k-1,k+1,\ldots,n\}$, we can rewrite
$\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}$ $=\big{(}\nabla^{2}f_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{k}\partial_{i}p_{t}(\mathbf{a}_{t})-\big{(}\nabla^{2}f_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{i}\partial_{k}p_{t}(\mathbf{a}_{t})$
$+\big{(}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{i}\partial_{k}f_{t}(\mathbf{a}_{t})-\big{(}\nabla^{2}p_{t}(\mathbf{a}_{t})\mathbf{v}\big{)}_{k}\partial_{i}f_{t}(\mathbf{a}_{t})$
$=\big{(}A(\mathbf{a}_{t},t)\mathbf{v}\big{)}_{k}\partial_{i}p_{t}(\mathbf{a}_{t})-\big{(}A(\mathbf{a}_{t},t)\mathbf{v}\big{)}_{i}\partial_{k}p_{t}(\mathbf{a}_{t})$
where we have used that $\partial_{i}f(\mathbf{a}_{t})=\lambda_{t}\partial_{i}p_{t}(\mathbf{a}_{t})$. By Lemma 4, $\nabla\mathcal{Q}_{k,i}(\mathbf{a}_{t},t)^{t}\mathbf{v}=0$ for all $i\in\{1,\ldots,n-1\}$ if and only if $A(\mathbf{a}_{t},t)\mathbf{v}$ is parallel to $\nabla p_{t}(\mathbf{a}_{t},t)$. Since we have
$\mathbf{v}^{t}A(\mathbf{a}_{t},t)\mathbf{v}<0\quad\text{ and }\mathbf{v}^{t}\nabla p_{t}(\mathbf{a}_{t})=0,$
this is a contradiction. ∎
Corollary 21.
Assume the same as in Theorem 6, and further that $f_{t}$ is concave, such that for some $k$ we have $\partial_{k}f_{t}\neq 0$ on $\mathbb{R}^{n}$ for every $t\in[0,1]$. Moreover, let $\mathbf{a}_{0}$ be the optimal value of
$\max_{\mathbf{x}\in C_{0}}\ f_{0}(\mathbf{x})$
Then the unique solution of (22) with initial value $\mathbf{a}_{0}$ for $t=0$ corresponds to the optimal values of the optimization problems (21).
Proof.
Identical to the proof of Theorem 6. ∎
Example 22.
Let
$C_{1}=\{\mathbf{x}\in\mathbb{R}^{n}\mid p_{1}(\mathbf{x})\geqslant 0\},$
where $p_{1}$ is a three times differentiable, concave function. In the following, we want to compute the closest point/distance of a point $\mathbf{y}\notin C_{1}$ to $C_{1}$. This corresponds to the following optimization problem:
$\max_{\mathbf{x}\in\mathbb{R}^{n}}$ $-\|\mathbf{x}-\mathbf{y}\|_{2}$
subject to $p_{1}(\mathbf{x})\geqslant 0.$
Let $p_{0}(\mathbf{x})=1-\|\mathbf{x}\|_{2}^{2}$. Moreover, let $f_{0}(\mathbf{x})=\mathbf{w}^{t}\mathbf{x}$ with $0\neq\mathbf{w}\in\mathbb{R}^{n}$ and ${f_{1}(\mathbf{x})=-\|\mathbf{x}-\mathbf{y}\|_{2}}$.
We consider the homotopy of objective functions
$f_{t}(\mathbf{x})=(1-t)f_{0}(\mathbf{x})+tf_{1}(\mathbf{x})$
Note that the optimal point of the problem for $t=0$ is given by $\mathbf{a}_{0}=\mathbf{w}/\|\mathbf{w}\|_{2}$. In Fig. 10 we show a numeric example for the path of optimal solutions for the concave function
$p_{1}(\mathbf{x})=5-\log\left(\sum_{k=1}^{4}\exp(\mathbf{b}_{k}^{t}\mathbf{x})\right),$
where $\mathbf{b}_{1}=(2,2)^{t}$, $\mathbf{b}_{2}=(-\frac{8}{5},\frac{8}{5})^{t}$, $\mathbf{b}_{3}=(-\frac{3}{5},\frac{6}{5})$, and $\mathbf{b}_{4}=(0,-\frac{6}{5})^{t}$. The inital linear function is given by $\mathbf{w}=(-1,1)^{t}$ and the distance is minimized to $\mathbf{y}=(-6,\frac{5}{2})^{t}\in C_{1}$.
(a) The initial optimization problem, i.e. optimizing $\mathbf{x}\mapsto x_{2}-x_{1}$ on a Euclidean ball.
(b) The target optimization problem. Computing the minimal distance of the point $(-6,2.5)$ to a convex set.
Figure 10: The homotopy method applied to the optimization problem described in Example 22. The path of optimal points is shown in both figures in grey.
Appendix C Boundedness of the Rigidly Convex Sets of Symmetric Polynomials

### v2

@@ 3.1.3 A Regular Homotopy for Real Zero Polynomials
Let $p_{0},p_{1}$ be real zero polynomials with bounded rigidly convex sets $\mathcal{R}(p_{0})$ and $\mathcal{R}(p_{1})$. Moreover, let $\varepsilon\colon[0,1]\to[0,\infty)$ be a smooth function with $\varepsilon(0)=\varepsilon(1)=0$ and $\varepsilon(t)>0$ for $t\in(0,1)$. We define the homotopies
@@
    $p_{t}^{[k]}(\mathbf{x})\coloneqq\mathcal{F}_{\varepsilon(t)}^{k}[\hat{p}_{t}](\mathbf{x}).$
Theorem 16.
@@
    If $p_{0}$ and $p_{1}$ are smooth, then $p_{t}^{[1]}$ is smooth for every $t\in[0,1)$.
  3. (iii)
    If $p_{0}$ is smooth and $p_{1}$ has roots with multiplicity $k\geq 2$, then $p_{t}^{[k]}$ is smooth for every $t\in[0,1)$.
@@
(iii) If $p_{0}$ and $p_{1}$ has roots with multiplicity $k$, we have that $s\mapsto\hat{p}_{t}(s\mathbf{x})$ has zeros of multiplicity at most $k+1,$ for every $t\in(0,1)$ and $0\neq\mathbf{x}\in\mathbb{R}^{n}$. Hence, applying the smoothing operator $\mathcal{F}_{\varepsilon(t)}^{k}$ gives rise to a smooth polynomial $p_{t}$ for every $t\in[0,1)$.
(ii) Special case of (iii) with $k=1$. ∎
Note that $\hat{p}_{t}$ is in general not smooth, even if $p_{0}$ and $p_{1}$ are smooth (see for example Fig. 2 for $\varepsilon=0$ and $t=0.6$). More specifically, the multiplicity of roots in $\hat{p}_{t}$ is in worst case the sum of the highest multiplicity arising in $p_{0}$ and $p_{1}$.
@@
If $p_{0}$ is smooth, $p_{1}$ has roots with multiplicity $k$, and $\mathcal{R}(p_{0}),\mathcal{R}(p_{1})$ are bounded, then the homotopy $p_{t}^{[k]}$ from Definition 15 is regular, and therefore the induced differential equation (6) has a unique solution.
@@
Regularity follows from smoothness of the $p_{t}^{[k]}$. Indeed, if a point leaves the innermost ring of zeros of $p_{t}$ over time, this would imply a multiple zero of some $p_{t}^{[k]}$ on some line. ∎
For non-smooth target sets, the smoothing degree $k$ must be sufficiently large to ensure the uniqueness of the solution in (6). However, in practice, the numerical solution obtained from the differential equation often converges to the correct result even for smaller values of $k$ (see Fig. 3). This behavior may be attributed to two factors:
  1. (i)
    The path of optimal points does not encounter any root of $\hat{p}_{t}$ with worst-case multiplicity, preserving the uniqueness of the solution.
  2. (ii)
    The numerical integrator, due to discretization in $t$, effectively bypasses the critical points.
Remark A similar statement toTheorem 16 can also be shown when dropping the smoothness-assumption for $\mathcal{R}(p_{1})$. In this situation $s\mapsto\hat{p}_{t}(s\mathbf{a})$ can have zeros up to multiplicity $d\coloneqq\deg(p_{1})$. Applying $\mathcal{F}_{\varepsilon(t)}$ $d$ times to $\hat{p},$ instead of only once, gives rise to a homotopy that satisfies the assumptions.
Figure 3: Optimizing the linear function $f(x_{1},x_{2})=x_{2}$ over a (non-smooth) spectrahedron (blue). The three paths are obtained by using the the homotopies $\hat{p}_{t}$ (left, blue), $p_{t}^{[1]}$ (middle, red) and $p_{t}^{[2]}$ (right, green). While only $p^{[2]}$ is a smooth homotopy, numerically all paths lead to a correct solution. The blue path on the left has two singularities, the red path in the middle has only one singularity and the green path on the right has no singularity.
@@ 4.1 Hyperbolic Programming
For this purpose, recall the notion of hyperbolic polynomials. Let ${h\colon\mathbb{R}^{n+1}\to\mathbb{R}}$ be a homogeneous polynomial. We call $h$ hyperbolic in the direction $\mathbf{e}\in\mathbb{R}^{n+1}$ if the univariate polynomial $t\mapsto h(\mathbf{y}-t\mathbf{e})$ has only real roots, for every $\mathbf{y}\in\mathbb{R}^{n+1}$. Hyperbolic polynomials give rise to a closed convex cone
@@
where $\mathbf{y}=(-\sum_{i=1}^{n}x_{i},x_{1},\ldots,x_{n})$. Moreover, the rigidly convex set $\mathcal{R}(p_{k})$ is bounded for $k\geqslant 2,$ see Appendix A. Note that $\mathcal{R}(p_{k})$ corresponds (up to an isometry) to the set
@@
Regarding (ii), we use the best known spectrahedral representation of $\mathcal{C}_{\mathbf{e}}(p_{k})$ introduced in [5]. According to the construction, the matrix size of this representation corresponds to the number of strings of length at most $k-1$ with distinct characters from an alphabet of size $n+1$. In other words, the matrix size $s(k,n)$ is given by
@@
$n=3$ $n=4$ $n=5$ $n=6$ $n=7$ $n=8$
$k=n-1$ $5$ $26$ $157$ $\mathbf{1100}$ $\mathbf{8801}$ $\mathbf{79210}$
$k=\lfloor\frac{n+1}{2}\rfloor$ $5$ $6$ $37$ $50$ $\mathbf{401}$ $\mathbf{5861}$
@@
4.2 Optimizing over the $k$-ellipsoids
@@
Note that $p$ is not differentiable at the focal points $\mathbf{u}_{i}$. Even if $\mathbf{u}_{i}\notin\partial\mathcal{E}_{n}(\mathbf{u}_{1},\ldots,\mathbf{u}_{k},r)$, there are times $t$ where the functions $p_{t}(\mathbf{x})$, constructed from the homotopy in Section 3.2, might not be differentiable and strictly concave. However, this happens for at most $k$ many different times. Therefore, the numerical integrator evaluates $p_{t}(\mathbf{x})$ on these problematic points with probability 00. At all other points and times, the function is concave in our above sense, so the method will work with probability $1$. The homotopy method for a $3$-ellipse is shown in Fig. 6(b).
@@
Note that (16) is a linear matrix inequality whose matrix size is exponential in the number of focal points $k$. There is also no significantly more efficient way of defining the $k$-ellipse by a linear matrix inequality, since the degree of the defining polynomial also increases exponentially with $k$. More precisely, this degree is $2^{k}$ if $k$ is odd and $2^{k}-\binom{k}{k/2}$ if $k$ is even [22]. Fig. 6(a) shows the zero set of the real zero polynomial $\det(\mathcal{M}(\mathbf{x}))$ for a $3$-ellipse.
(a) The zero set of the real zero polynomial $\det(\mathcal{M}(\mathbf{x}))$. The loop surrounding $\mathbf{0}$ bounds the $3$-ellipse.
(b) The homotopy method applied to the $3$-ellipse and the functional ${f(\mathbf{x})=x_{1}+x_{2}}$. The figure shows the sets at time 00, $0.7$, $0.9$ and $1,$ and the path of optimal solutions.
@@
$k=2$ $k=3$ $k=4$ $k=5$ $k=10$ $k=100$
$n=2$ Homotopy $0.02(1)$ $0.02(1)$ $0.06(4)$ $0.04(1)$ $0.09(4)$ $1.5(4)$
SDP $0.03(1)$ $0.07(1)$ $3.3(4)$ $\approx 1500$* - -
$2^{\text{nd}}$ order cone $0.02(1)$ $0.02(1)$ $0.03(1)$ $0.03(2)$ $0.04(2)$ $0.15(3)$
$n=4$ Homotopy $0.05(2)$ $0.05(1)$ $0.07(3)$ $0.10(2)$ $0.21(5)$ $4(1)$
SDP $0.06(1)$ $23(2)$
$2^{\text{nd}}$ order cone $0.02(1)$ $0.03(1)$ $0.03(1)$ $0.03(1)$ $0.05(2)$ $0.22(3)$
@@ 4.3 Optimizing over a $p$-Norm Ball
For the SDP representation, note that $B_{8,r}$ is not a spectrahedron (see [20, Example 2.29] for a similar argument for $B_{4,r}$). However, it can be represented as the projection of a spectrahedron. This leads to the following SDP which is equivalent to (18) for $p=8$.
@@ 4.4 Geometric Programming with a Single Constraint
(a) Varying the number of monomials $r$ for fixed dimension $n=4$.
(b) Varying the dimension $n$. For the number of monomials we choose $r=4n$ to guarantee that the feasable set is bounded with high probability.
@@ 5 Conclusion and Open Questions
It also remains unclear if the boundedness requirement of the feasible sets $C_{t}$ can be relaxed. While boundedness limits hyperbolic programming only to affine slices of hyperbolicity cones, it is crucial for preventing solution divergence, even though smoothing also works in the unbounded case.
Moreover, we leave an exact runtime analysis of this method for further work. This would include for example a comparison of the complexity of our method with interior-point methods applied to SDPs. Finally, since there are multiple canonical choices of homotopies for certain optimization classes this suggests to study whether there exist alternative choices of homotopies that lead to computationally more efficient results.
@@ Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) [doi:10.55776/P33122]. AK further acknowledges funding of the Austrian Academy of Sciences (ÖAW) through the DOC scholarship 26547. Moreover, AK wants to thank Hamza Fawzi, and Oisín Faust for fruitful discussions.
Appendix A Boundedness of the Rigidly Convex Sets of Symmetric Polynomials